find the value of r so the line that passes through each pair of the points has the given slope.
1) (1,4) (r,5), m= undefined

Answer by stanbon(75887)

find the value of r so the line that passes through each pair of the points has the given slope.
1) (1,4) (r,5), m= undefined
---------------
slope = (5-4)/(r-1) = undefined
r-1 = 0
r = 1
===========
Cheers,
Stan H.